# Online Feature Updates Improve Online (Generalized) Label Shift Adaptation

### v1

@@ 3 Method
We introduce a novel online label shift adaptation algorithm OLS-OFU, that leverages self-supervised learning (SSL) to improve representation learning, while maintaining theoretical guarantees. This approach is general and can be seamlessly integrated with any existing online label shift (OLS) method. We show both theoretical and empirical improvement of OLS-OFU. Moreover, we demonstrate their effectiveness in addressing the more challenging problem of online generalized label shift adaptation.
3.1 Online Label Shift Adaptation with Online Feature Updates
Main ideas.
Before formally introducing the algorithm, we would like to discuss two crucial design principles about where to place the update feature extractor through SSL, inspired by the analysis of OLS methods in the literature. These principles are essential for maintaining both the theoretical guarantees and potential empirical performance .
Principle 1: at time step $t$, the feature extractor update using samples $S_{t}$ should occur after the OLS updates.
The main idea of ROGD, UOGD, and ATLAS is to construct an unbiased estimator for the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$ using a batch of samples $S_{t}\sim\mathcal{P}_{t}^{\rm test}$. The estimator has the form of
$\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{x\sim\mathcal{P}^{\rm train}(\cdot|y)}\ell_{\rm sup}(f_{t}(x),y),$ (3)
where the randomness of $s_{t}$ depends on samples $S_{t}$ as constructed and is an unbiased estimator for the label marginal distribution $q_{t}$. The following proposition introduces a necessary condition of $f_{t}$ to ensure the unbiased property of the above estimator and the proof is given at Appendix B.
@@ 3.1 Online Label Shift Adaptation with Online Feature Updates
From the proposition, it is necessary to ensure $f_{t}$ is independent of $S_{t}$. Thus, the update of the feature extractor in $f_{t}$ (using SSL) using samples $S_{t}$ should occur after the online label shift (OLS) method is updated with $S_{t}$.
Principle 2: the last linear layer should be re-trained using the train set $D_{0}$ after the feature extractor update.
The key observation here is that a label-shifted distribution $\mathcal{P}^{\rm test}_{t}(y|x)$ is an exact post-hoc reweight of $\mathcal{P}^{\rm train}(y|x)$:
$\mathcal{P}^{\rm test}_{t}(y|x)=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\mathcal{P}^{\rm test}_{t}(x|y)=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\mathcal{P}^{\rm train}(x|y)$
$=\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm test}_{t}(x)}\frac{\mathcal{P}^{\rm train}(x)}{\mathcal{P}^{\rm train}(y)}\mathcal{P}^{\rm train}(y|x)\propto\frac{\mathcal{P}^{\rm test}_{t}(y)}{\mathcal{P}^{\rm train}(y)}\mathcal{P}^{\rm train}(y|x).$
Another observation is that $f_{0}$ trained from $D_{0}$ could be a good approximation of $\mathcal{P}^{\rm train}(y|x)$. These two observations motivate approaches like ROGD, FTH, and FLHFTL to opt for the hypothesis $f$ to be a post-hoc reweight of the pre-trained model $f_{0}$. In the context of our framework, these methods need to be adapted slightly. Instead of reweighting the original model $f_{0}$ directly, they should reweight a model within the updated feature extractor. This reweighted model should still approximate the original train data distribution $\mathcal{P}^{\rm train}(y|x)$, which can be achieved by re-training the last linear layer using the initial train data $D_{0}$.
Algorithm details.
@@
The two principles together motivate our final algorithm: Online Label Shift adaptation with Online Feature Updates (OLS-OFU; Algorithm 1), which requires a self-supervised learning loss $\ell_{\rm ssl}$ and an online label shift adaptation algorithm (OLS) that either reweights the offline pre-trained model $f_{0}$ or updates the last linear layer [^1] As pointed out in Section 2, this is relatively general as most previous online label shift algorithms belong to these two categories.. In the training stage, we train $f_{0}$ by minimizing the supervised and self-supervised loss together defined on train data. In the test stage, OLS-OFU comprises three steps at each time step $t$: (1) running the refined version of OLS, which we refer to as OLS-R, (2) updating the feature extractor, and (3) re-training the last linear layer. The details of these three steps are elaborated below.
@@
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, train set $D_{0}$ and validation set $D_{0}^{\prime}$.
     1. Run the revised version of OLS, that is, OLS-R, and get $f_{t+1}^{\prime}$. (Algorithm 2 is the revised FLHFTL.)
@@
     3. Re-train the last linear layer by Equation 5. Calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration on $D_{0}^{\prime}$ and get the calibrated model $f_{t+1}^{\prime\prime}$. Output at time $t$: For the reweighting OLS methods, denote the latest reweighting vector from Step 1 is $p_{t+1}$ and we define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$; else, we define $f_{t+1}:=f_{t+1}^{\prime}$.
@@
Algorithm 2 Revised FLHFTL for online feature updates (FLHFTL-R); See the original version in (Baby et al., 2023).
  Require: Online regression oracle ALG.
  for $t=1,\cdots,T$ do
[…]
     2. Compute $\tilde{q}_{t+1}:={\rm ALG}(s_{1},\cdots,s_{t})$
[…]
     Output at time $t$: $f_{t+1}^{\prime}$.
  end for
(1) Running the Revised OLS. At the beginning of the time $t$, we use $f_{t}^{\prime\prime}$ to denote the model within the latest feature extractor and the re-trained linear model (using data $\{S_{0},\cdots,S_{t-1}\}$). The high level idea of our framework centers on substituting the pre-trained model $f_{0}$ used in existing OLS methods with our updated model $f_{t}^{\prime\prime}$, which we call Revised OLS. To provide an overview, we examine common OLS methods (FLHFTL, FTH, ROGD, UOGD, and ATLAS) and identify two primary usecases of $f_{0}$. Firstly, all OLS methods rely on an unbiased estimator $s_{t}$ of the label distribution $q_{t}$ and $f_{0}$ is a part of the estimator. Secondly, the hypothesis $f$ is either reweighted based on $f_{0}$ or only the last linear layer in $f_{0}$ is updated. To illustrate the revised OLS formally, Algorithm  2 showcases the specific modifications applied to FLHFTL as an example. Further details regarding the revisions for ROGD, FTH, UOGD, and ATLAS can be found in Appendix C. We use $f_{t+1}^{\prime}$ to denote the model after running the revised OLS.
(2) Updating the Feature Extractor. We now introduce how to utilize the SSL loss $\ell_{\rm ssl}$ to update the feature extractor based on any incoming batch of unlabeled test data $S_{t}$ at timestep $t$. Specifically, let $\theta^{\rm feat}_{t}$ denote the parameters of the feature extractor in $f_{t+1}^{\prime}$. The update of $\theta^{\rm feat}_{t+1}$ at time $t$ is given by:
$\theta_{t+1}^{\rm feat}:=\theta_{t}^{\rm feat}-\eta\cdot\nabla_{\theta^{\rm feat}}\ell_{\rm ssl}(S_{t};f_{t+1}^{\prime}).$ (4)
@@ 2. Compute $\tilde{q}_{t+1}:={\rm ALG}(s_{1},\cdots,s_{t})$
(3) Re-training Last Linear Layer. With the updated feature extractor $\theta_{t+1}^{\rm feat}$, it is important to re-train the last linear layer $\theta^{\rm linear}_{t+1}$ to align with the new feature extractor. The re-training starts from random initialization, while the feature extractor remains frozen. The training objective of $\theta^{\rm linear}_{t+1}$ is to minimize the average cross-entropy loss over train data $D_{0}$. We denote the model with the frozen feature extractor $\theta_{t+1}^{\rm feat}$ as $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear})$. The objective for re-training last linear layer can be written as follows:
@@
We calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration (Guo et al., 2017) using the validation set $D_{0}^{\prime}$ and denote the model after calibration as $f_{t+1}^{\prime\prime}$. This re-training step is essential to ensure the model is a calibrated model of estimating $\mathcal{P}^{\rm train}(y|x)$ rather than other distributions.
Ouput at time $t$. Finally, we define $f_{t+1}$ for the next time step prediction. For the reweighting OLS methods (ROGD, FTH, FLHFTL), denote the latest reweighting vector from Step 1 is $p_{t+1}$ and define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$. For those which optimize the last linear layer (UOGD, ATLAS), we define $f_{t+1}:=f_{t+1}^{\prime}$ – as stated in Principle 2 in Section 3.1, the $f_{t}^{\prime\prime}$ including retrained last linear layer serves specially for the reweighting OLS method.
Batch accumulation for SSL that needs large batches. The loss of SSL such as contrastive learning (Chen et al., 2020a; He et al., 2020) is defined from a batch of data and the effective batch is necessary to be large. We introduce a simple variant of Algorithm 1: we process Step 2-3 every BA time steps by the samples collected in the period.
3.2 Performance Guarantee
The original OLS methods exhibit theoretical guarantees in terms of regret convergence for online label shift setting, where $\mathcal{P}^{\rm test}_{t}(x|y)=\mathcal{P}^{\rm train}(x|y)$ is invariant. With the incorporation of the additional online feature update step, OLS-OFU demonstrates analogous theoretical results. By comparing the theoretical results between OLS and OLS-OFU, we can gain insights into potential enhancements from OLS to OLS-OFU. Due to limited space, we provide the theoretical results pertaining to FLHFTL-OFU here, and present the theoretical results for ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU in Appendix D.
@@ 3.2 Performance Guarantee
[Regret convergence for FLHFTL-OFU] Suppose we choose the OLS subroutine in Algorithm 2 to be FLHFTL from Baby et al. (2023). Let $f_{t}^{\rm flhftl-ofu}$ be the output at time step $t-1$ from Algorithm 1, that is $g(\cdot;f_{t}^{\prime\prime},\frac{\tilde{q}_{t}}{q_{0}})$. Let $\sigma$ be the smallest among the the minimum singular values of invertible confusion matrices $\{C_{f_{1}^{\prime\prime},D_{0}^{\prime}},\cdots C_{f_{T}^{\prime\prime},D_{0}^{\prime}}\}$. Then under Assumptions 1 and 2 in Baby et al. (2023), FLHFTL-OFU has the guarantee below: $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]$ $\leq O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (6)
@@
To ease comparison, we state the guarantee for the original OLS method FLHFTL (Baby et al., 2023):
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl};\mathcal{P}_{t}^{\rm test})\right]-\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{0},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})$ $\leq O\left(\frac{K^{1/6}V_{T}^{1/3}}{\sigma^{2/3}T^{1/3}}+\frac{K}{\sigma\sqrt{T}}\right),$ (7)
Recall that the objective function for the online label shift problem is defined as the average loss in Equation 1. Both theorems establish the convergence in terms of this average loss. In the event that $f_{t}^{\prime\prime}$ ($t\in[T]$) from the online feature updates yield improvements:
@@
then it guarantees that the loss of FLHFTL-OFU will converge to a smaller value, resulting in enhanced performance compared to FLHFTL. We substantiate this improvement through empirical evaluation in Section 4. For other OLS methods such as ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU, a similar analysis can be derived and we present them in Appendix D.

### v3

@@ 3 Method
In this paper, we introduce a novel online label shift adaptation algorithm OLS-OFU, that leverages self-supervised learning (SSL) to improve representation learning and can be seamlessly integrated with any existing online label shift (OLS) method. By carefully designing how to place the SSL techniques, our algorithm maintains a similar theoretical guarantee from the existing OLS methods, obeys the underlying assumption in existing OLS methods that is important for the effectiveness, and fits the required condition of the SSL techniques while avoiding heavy additional time cost. Through the derived theoretical results, we further understand how the online learning techniques and SSL techniques contribute together to reduce the test loss in the online label shift adaptation problem. Lastly, we demonstrate how the SSL techniques help with the more challenging problem online generalized label shift adaptation.
3.1 Three Principles of Algorithm Design.
To combine the step of feature extractor update with any OLS method, the most straightforward thought is to directly insert this step right before or after the step of original OLS in each time step; see the summarization of the online process in Figure 1. Starting from this thought, there are three remaining questions before finalizing the algorithm: Q1: Inserting this step before or after the original OLS step, which option is better? Q2: Besides this feature extractor update step, are any other steps also necessary? Q3: How frequently (in terms of time step $t$) should we update the feature extractor? We are going to introduce three principles when designing our algorithm, which helps answer the three questions and together lead to our final design.
Principle 1: maintain the theoretical guarantee. One main advantage of existing OLS methods is that they have theoretical guarantees for the performance of any unknown label shifts in the online test stage. We would like to keep this advantage after deploying the feature extractor update step. By carefully checking the theoretical analysis of the existing OLS methods, we find that when we deploy the feature extractor update into ROGD, UOGD, or ATLAS, if we would like to update the feature extractor by the unlabeled test samples including $S_{t}$, it is necessary to execute this update after the step of these OLS methods at time $t$ to maintain the similar theoretical guarantee. This is because the main idea of ROGD, UOGD, and ATLAS is to construct an unbiased estimator for the gradient $\nabla_{f}\ell(f_{t};\mathcal{P})$ using a batch of samples $S_{t}\sim\mathcal{P}_{t}^{\rm test}$. The estimator has the form of
$\sum_{y\in\mathcal{Y}}s_{t}[y]\cdot\nabla_{f}\mathbb{E}_{(x,y)\sim\mathcal{P}^{\rm train}}\ell_{\rm sup}(f_{t}(x),y),$ (3)
where $s_{t}$ is an unbiased estimator for the label marginal distribution $q_{t}$ and it depends on samples $S_{t}$ as constructed. If the feature extractor update according to $S_{t}$ happened before the step of ROGD, UOGD or ATLAS, $f_{t}$ would not be independent of $s_{t}$ in Equation 3 and this can break their main idea, as illustrated in the following proposition.
@@
Thus, to maintain a similar theoretical guarantee after updating the feature extractor by the online test samples, we should insert the feature update involving $S_{t}$ after the step of original ROGD, UOGD or ATLAS, and this answers Q1. In the experiment, we will validate the necessity of this design.
Principle 2: obey the underlying assumption of the existing OLS methods. We find that the existing OLS methods ROGD, FTH, and FLHFTL opt for the hypothesis $f$ to be a post-hoc reweight of the pre-trained model $f_{0}$. The underlying assumption behind this design is that $f_{0}$ is a good approximation of $\mathcal{P}^{\rm train}(y|x)$. Within this assumption, because $\mathcal{P}^{\rm test}(y|x)$ is a post-hoc reweight of $\mathcal{P}^{\rm train}(y|x)$ in the case of label shift [34, 49], reweighting such $f_{0}$ can approach $\mathcal{P}^{\rm test}(y|x)$. Therefore, after updating the feature extractor, we expect that the base model, which is to be reweighted, still approximates the training distribution $\mathcal{P}^{\rm train}(y|x)$. This can be done by re-training the linear layer under the training data after the feature extractor update, and this answers Q2.
Principle 3: fit the required condition of the SSL techniques while avoiding heavy additional computational costs. Given a set of unlabeled samples $S$, we denote the loss of an SSL technique for a model $f$ as $\ell_{\rm ssl}(S;f)$ and the update to this model would be in the form of gradient descent: $\theta^{\rm feat}\to\theta^{\rm feat}-\eta\cdot\nabla_{\theta^{\rm feat}}\ell_{\rm ssl}(S;f)$, where $\eta$ is the learning rate. We notice that the batch size $|S|$ cannot be very small. Otherwise, the update can be too noisy due to the data variance or some types of SSL whose benefit replies on large batch sizes, such as contrastive learning [9, 23], would lose this benefit. However, online (generalized) label shift problem assumes we only have a small batch of unlabeled samples $S_{t}$ at time $t$; the batch size is set as $1$ or $10$ in the experiment of literature. Therefore, for the effectiveness of SSL techniques, instead of updating the feature extractor at each time step by $S_{t}$, we opt to accumulate the sample batch $S_{\tau c}\cup\cdots\cup S_{\tau(c+1)}$ and update the feature extractor once every $\tau$ time steps, and we call it batch accumulation.
Besides the help of effectiveness, batch accumulation also helps with time efficiency. The existing OLS methods only involve the updates for the small reweighting vector or the linear layer and the time cost is low. However, updating the feature extractor and re-training the linear layer on the full training set, which are introduced in Principle 2, are much computationally heavier. For example, when we evaluated the methods in the experiment with ResNet18 on CIFAR10 dataset, one step of FLHFTL only took $0.069$ second, while one step of FLHFTL together with feature extractor update and linear layer re-training took $17.1$ seconds, which is $247$ times. After applying batch accumulation and we select $\tau=100$, the additional time costs will be reduced by $(\tau-1)/\tau=99\%$.
Batch accumulation answers $Q_{3}$we should update the feature extractor every $\tau$ steps for effectiveness and time efficiency. We will study how batch accumulation helps in the experiments.
3.2 Online Label Shift Adaptation with Online Feature Updates
@@
We have discussed three principles and now we can finalize our algorithm Online Label Shift adaptation with Online Feature Updates (OLS-OFU; Algorithm 1), which requires a self-supervised learning loss $\ell_{\rm ssl}$ and one of existing OLS methods in the literature, which either reweights the offline pre-trained model $f_{0}$ or updates the last linear layer. In the training stage, we train $f_{0}$ by minimizing the supervised and self-supervised loss together defined on train data. In the test stage, OLS-OFU comprises three steps at each time step $t$: (1) running the refined version of OLS, which we refer to as OLS-R, (2) updating the feature extractor, and (3) re-training the last linear layer. As suggested by Principle 3 in Section 3.1, steps (2-3) only run every $\tau$ steps. We illustrate the details of these three steps are elaborated below.
@@
     Input at time $t$: Samples $S_{1}\cup\cdots\cup S_{t}$, models $\{f_{1},\cdots,f_{t}\}$, train set $D_{0}$, validation set $D_{0}^{\prime}$.
     1. Run the revised version of OLS, that is, OLS-R, and get $f_{t+1}^{\prime}$.
     If $t\%\tau\neq 0$, $f_{t+1}:=f_{t+1}^{\prime}$, $f_{t+1}^{\prime\prime}:=f_{t}^{\prime\prime}$ Else, go to the next step 2-3.
@@
         3. Re-train the last linear layer by Equation 5, calibrate the model and get $f_{t+1}^{\prime\prime}$. Output at time $t$: For the reweighting OLS methods, denote the latest reweighting vector from Step 1 is $p_{t+1}$ and we define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$; else, we define $f_{t+1}:=f_{t+1}^{\prime}$.
@@
(1) Running the Revised OLS. At the beginning of the time $t$, we use $f_{t}^{\prime\prime}$ to denote the model within the latest feature extractor and the re-trained linear model (using data $\{S_{0},\cdots,S_{t-1}\}$). The high level idea of our framework centers on substituting the pre-trained model $f_{0}$ used in existing OLS methods with our updated model $f_{t}^{\prime\prime}$, which we call OLS-R. To provide an overview, we examine common OLS methods (FLHFTL, FTH, ROGD, UOGD, and ATLAS) and identify two primary use cases of $f_{0}$. Firstly, all OLS methods rely on an unbiased estimator $s_{t}$ of the label distribution $q_{t}$ and $f_{0}$ is a part of the estimator. Secondly, the hypothesis $f$ is some weights of reweighting the $f_{0}$ output or only the last linear layer in $f_{0}$ is updated. We illustrate all revised OLS methods formally in Appendix C. We use $f_{t+1}^{\prime}$ to denote the model after running the OLS-R.
(2) Updating the Feature Extractor. As guided by Principle 1 in Section 3.1, the step of updating the feature extractor should be after the Revised-OLS module (step (1)) for theoretical guarantees. We now introduce how to utilize the SSL loss $\ell_{\rm ssl}$ to update the feature extractor based on the accumulated batch of unlabeled test data $S_{\tau c}\cup\cdots\cup S_{\tau(c+1)}$ at timestep $t=\tau(c+1)$. Specifically, let $\theta^{\rm feat}_{t}$ denote the parameters of the feature extractor in $f_{t+1}^{\prime}$. The update of $\theta^{\rm feat}_{t+1}$ is given by:
$\theta_{t+1}^{\rm feat}:=\theta_{t}^{\rm feat}-\eta\cdot\nabla_{\theta^{\rm feat}}\ell_{\rm ssl}(S_{\tau c}\cup\cdots\cup S_{\tau(c+1)};f_{t+1}^{\prime}).$ (4)
@@
(3) Re-training Last Linear Layer. Principle 2 in Section 3.1 suggests this step of re-training the last linear layer on the training distribution. The re-training starts from random initialization while the feature extractor remains frozen. The training objective of $\theta^{\rm linear}_{t+1}$ is to minimize the average cross-entropy loss over train data $D_{0}$. We denote the model with the frozen feature extractor $\theta_{t+1}^{\rm feat}$ as $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear})$. The objective for re-training the last linear layer can be written as follows:
@@
We calibrate the model $f(\cdot|\theta^{\rm feat}_{t+1},\theta^{\rm linear}_{t+1})$ by temperature calibration [21] using the validation set $D_{0}^{\prime}$ and denote the model after calibration as $f_{t+1}^{\prime\prime}$.
Output at time $t$. Finally, we define $f_{t+1}$ for the next time step prediction. For the reweighting OLS methods (ROGD, FTH, FLHFTL), denote the latest reweighting vector from Step 1 as $p_{t+1}$ and define $f_{t+1}:=g(\cdot;f_{t+1}^{\prime\prime},p_{t+1})$. For those which optimize the last linear layer (UOGD, ATLAS), we define $f_{t+1}:=f_{t+1}^{\prime}$ – as stated in Principle 2 in Section 3.2, the $f_{t}^{\prime\prime}$ including retrained last linear layer serves specially for the reweighting OLS method.
The requirement of the training data. Notice that in step (3), we need the training data to retrain the linear classifier, which potentially brings the additional cost of memory and computation in the test time. As discussed in Principle 2 and lines 248-252, retraining the last linear layer is designed only for three previous OLS methods (ROGD, FTH, FLHFTL); our algorithm for two other OLS methods (UOGD and ATLAS) in the literature are independent of this step. As for ROGD, FTH, or FLHFTL, we interestingly find that OLS-OFU without this re-training step, which means that we update the feature extractor but reuse the pretrained linear classifier, still has a certain advantage over the original OLS methods – although more stored training data results in the more significant benefit of OLS-OFU. This suggests that in practice, if we reduce the amount of stored training data due to the memory or computational constraint, OLS-OFU is still effective.
3.3 Performance Guarantee for Online Label Shift Adaptation
One main advantage of the original OLS methods is that they exhibit theoretical guarantees in terms of regret convergence for online label shift settings. Next, we will show how OLS-OFU demonstrates analogous theoretical guarantees with the incorporation of additional online feature updates. Due to limited space, we illustrate the theoretical results pertaining to FLHFTL-OFU here, and present the results for ROGD-OFU, FTH-OFU, UOGD-OFU, and ATLAS-OFU in Appendix D.
@@ 3.3 Performance Guarantee for Online Label Shift Adaptation
[Regret convergence for FLHFTL-OFU] Suppose we choose the OLS-R to be FLHFTL-R (Algorithm 6) from Baby et al. [6]. Let $f_{t}^{\rm flhftl-ofu}$ be the output at time step $t-1$ from Algorithm 1, that is $g(\cdot;f_{t}^{\prime\prime},\frac{\tilde{q}_{t}}{q_{0}})$. Under Assumptions 1 and 2 in Baby et al. [6], FLHFTL-OFU has the guarantee:
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}^{\rm flhftl-ofu};\mathcal{P}_{t}^{\rm test})\right]\leq\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\ell(g(\cdot;f_{t}^{\prime\prime},\frac{q_{t}}{q_{0}});\mathcal{P}_{t}^{\rm test})\right]+O\left(\frac{K^{1/6}V_{T}^{1/3}}{T^{1/3}}+\frac{K}{\sqrt{T}}\right),$ (6)
@@
How do online feature updates contribute to the bound? We observe that the upper bound of the test loss has two terms. The first term is the loss of the model within the up-to-date feature extractor when the knowledge of label distribution $q_{t}$ is known. Any improvement from SSL could be reflected in this first term. The second term is to quantify the loss gap between the knowledge of label distribution $q_{t}$ and the estimation of the label distribution by the online learning technique.
When do online feature updates improve the guarantee from FLHFTL to FLHFTL-OFU? If we do not make any update to the feature extractor (i.e. $f_{t}^{\prime\prime}=f_{0},\forall t$), the upper bound in Equation 6 would be naturally reduced to the theoretical guarantee of FLHFTL [6]. Moreover, in the following event of $f_{t}^{\prime\prime}$ ($t\in[T]$):
@@
our theorem will guarantee that the loss of FLHFTL-OFU converges to a smaller value than the one of the original FLHFTL, resulting in a better upper bound. We substantiate this improvement through empirical evaluation in Section 4.